A body of mass 1.0kg is falling with an acceleration of 10 m\sec​2. Its apparent weight will be (g=10m\sec2)

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
R=m(g−a)=m(10−10)=zero
 
hence option D is correct
in case of qny query please feel free to ask
 
Regards
Arun (askIITians forum expert)